OpenStudy (anonymous):

if l ak l (k is as a subscript)<3 ,1

OpenStudy (anonymous):

\(|a_k|<3,1<k<n\) and consider all \(z\) satisfy \(0=1+a_1 z+a_2 z^2+a_3z^3+\dots+a_n z^n\)... so$$0=1+a_1z+\dots+a_nz^n\\0=|1+a_1z+\dots+a_n z^n|\le 1+|a_1||z|+\dots+|a_n||z|^n$$

OpenStudy (anonymous):

and $$1+|a_1||z|+\dots+|a_n||z|^n<1+3|z|+\dots+3|z|^n$$but$$1+3|z|+\dots+3|z|^n=1+3|z|(1+\dots+|z|^{n-1})=1+3|z|\cdot\frac{1-|z|^n}{1-|z|}$$

OpenStudy (anonymous):

hmm...

OpenStudy (anonymous):

http://en.wikipedia.org/wiki/Properties_of_polynomial_roots#Based_on_the_Rouch.C3.A9_theorem can't complete it but this is what they want you to do ^^ sorry I have to get ready for class